# Magnetic Field Configurations in Binary Neutron Star Mergers I: Post-merger Remnant and Disk

### v1

@@ IV Discussion
The toroidal field dominates over the poloidal field independently of the setup and this persists during the simulations (see Figs. 11 and the 4th row of Fig. 16), with the toroidal-to-poloidal energy ratio stabilizing around $2{-}4$ in most cases within the disk but with a much larger deviation (${\sim}8-30$) within the remnant. This ratio is particularly important for understanding the potential launching of relativistic jets, as the field configuration strongly influences the efficiency of energy extraction mechanisms. The magnetic-to-gas pressure ratio $\beta^{-1}$ shows a secular growth in our simulations, particularly in the disk (Fig. 12). It does not reach saturation within the simulations’ lifespan, suggesting that longer-term evolution might lead to even higher magnetization, potentially reaching the values necessary for jet launching at later times. The radial profiles shown in Figs. 15 and 16 reveal that $\beta^{-1}$ grows faster with time at larger radii, indicating that the disk becomes increasingly magnetically dominated over time. The growing magnetization in the disk may provide favorable conditions for launching outflows, though they are not sufficient yet to show a clear difference in the amount of ejected mass (see Paper II). The strong toroidal field in the remnant could support the formation of a central engine for sGRBs.
As shown in Fig. 3, simulations with initially vertical and aligned dipolar fields develop toroidal fields with a counter-clockwise (clockwise) direction above (below) the equatorial plane. In contrast, simulations with initially anti-aligned poloidal field (POL_UD) or a toroidal field (TOR_PP) give rise to toroidal fields which exhibit the opposite polarity. Despite these differences in field topology, the overall energetics of the system –particularly the toroidal-to-poloidal energy ratio and the $\beta^{-1}$ parameter– show consistency across different initial configurations. Interestingly, the simulation with bitant symmetry (BITANT) shows a significantly more coherent field structure on the equatorial plane even at early times after the merger, indicating that the imposed symmetry condition reduces the development of turbulence. This has methodological implications for future simulations, as symmetry assumptions may artificially suppress certain aspects of the magnetic field evolution.
@@ V Conclusions and future directions
In all the simulations, the mass of the disk rapidly grows at the merger and then saturates or grows at a low rate. Simulation POL_q=1.2 achieves the largest disk mass, ${\sim}0.14\leavevmode\nobreak\ M_{\odot}$, due to the tidal disruption of the lighter star before the merger, followed by STIFF, which, due to the larger volume of the remnant, allows more material to be loosely bound to the remnant and populate the disk. The rest of the simulations show consistent disk mass values of $\lesssim 0.1\leavevmode\nobreak\ M_{\odot}$. The influence of the magnetic field on the disk mass is unclear, since whereas LOW_B shows a larger mass than the other runs, simulation NO_B behaves similarly to the high-$B$ runs.
@@ Acknowledgements.
EG thanks Luciano Combi for helpful discussions. EG and DR acknowledge funding from the National Science Foundation under Grants No. AST-2108467 and PHY-2407681. EG acknowledges funding from an Institute for Gravitation and Cosmology fellowship. JF and DR acknowledge U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Number(s) DE-SC0021177 and DE-SC0024388. BD and SB acknowledge support by the EU Horizon under ERC Consolidator Grant, no. InspiReM-101043372. PH acknowledges funding from the National Science Foundation under Grant No. PHY-2116686. The authors are indebted to A. Celentano’s PRİSENCÓLİNENSİNÁİNCIÚSOL. The numerical simulations were performed on TACC’s Frontera (NSF LRAC allocation PHY23001) and on Perlmutter. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Simulations were also performed on the national HPE Apollo Hawk at the High Performance Computing Center Stuttgart (HLRS). The authors acknowledge HLRS for funding this project by providing access to the supercomputer HPE Apollo Hawk under the grant numbers INTRHYGUE/44215 and MAGNETIST/44288. Simulations were also performed on SuperMUC_NG at the Leibniz-Rechenzentrum (LRZ) Munich. The authors acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at LRZ (allocations pn68wi and pn36jo). Postprocessing and development run were performed on the ARA cluster at Friedrich Schiller University Jena. The ARA cluster is funded in part by DFG grants INST 275/334-1 FUGG and INST 275/363-1 FUGG, and ERC Starting Grant, grant agreement no. BinGraSp-714626.

### v2

@@ IV Discussion
The toroidal field dominates over the poloidal field independently of the setup and this persists during the simulations (see Figs. 12 and the 4th row of Fig. 17), with the toroidal-to-poloidal energy ratio stabilizing around 242{-}4 in most cases within the disk but with a much larger deviation (${\sim}8-30$) within the remnant. This ratio is particularly important for understanding the potential launching of relativistic jets, as the field configuration strongly influences the efficiency of energy extraction mechanisms. The magnetic-to-gas pressure ratio $\beta^{-1}$ shows a secular growth in our simulations, particularly in the disk (Fig. 13). It does not reach saturation within the simulations’ lifespan, suggesting that longer-term evolution might lead to even higher magnetization, potentially reaching the values necessary for jet launching at later times. The radial profiles shown in Figs. 16 and 17 reveal that $\beta^{-1}$ grows faster with time at larger radii, indicating that the disk becomes increasingly magnetically dominated over time. The growing magnetization in the disk may provide favorable conditions for launching outflows, though they are not sufficient yet to show a clear difference in the amount of ejected mass (see Paper II). The strong toroidal field in the remnant could support the formation of a central engine for sGRBs.
As shown in Fig. 3, simulations with initially vertical and aligned dipolar fields develop toroidal fields with a counter-clockwise (clockwise) direction above (below) the equatorial plane. In contrast, simulations with initially anti-aligned poloidal field (POL-UD) or a toroidal field (TOR-PP) give rise to toroidal fields which exhibit the opposite polarity. Despite these differences in field topology, the overall energetics of the system –particularly the toroidal-to-poloidal energy ratio and the $\beta^{-1}$ parameter– show consistency across different initial configurations. Interestingly, the simulation with bitant symmetry (BITANT) shows a significantly more coherent field structure on the equatorial plane even at early times after the merger, indicating that the imposed symmetry condition reduces the development of turbulence. This has methodological implications for future simulations, as symmetry assumptions may artificially suppress certain aspects of the magnetic field evolution.
@@ V Conclusions and future directions
In all the simulations, the mass of the disk rapidly grows at the merger and then saturates or grows at a low rate. Simulation POL-Q12 achieves the largest disk mass, ${\sim}0.14\leavevmode\nobreak\ M_{\odot}$, due to the tidal disruption of the lighter star before the merger, followed by STIFF, which, due to the larger volume of the remnant, allows more material to be loosely bound to the remnant and populate the disk. The rest of the simulations show consistent disk mass values of $\lesssim 0.1\leavevmode\nobreak\ M_{\odot}$. The influence of the magnetic field on the disk mass is unclear, since whereas LOWB shows a larger mass than the other runs, simulation B0 behaves similarly to the high-$B$ runs.
@@ Acknowledgements.
We thank the referee for the useful comments. EG thanks Luciano Combi for helpful discussions. EG and DR acknowledge funding from the National Science Foundation under Grants No. AST-2108467 and PHY-2407681. EG acknowledges funding from an Institute for Gravitation and Cosmology fellowship. JF and DR acknowledge U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Number(s) DE-SC0021177 and DE-SC0024388. BD, SB, and MJ acknowledge support by the EU Horizon under ERC Consolidator Grant, no. InspiReM-101043372. PH acknowledges funding from the National Science Foundation under Grant No. PHY-2116686. The authors are indebted to A. Celentano’s PRİSENCÓLİNENSİNÁİNCIÚSOL. The numerical simulations were performed on TACC’s Frontera (NSF LRAC allocation PHY23001) and on Perlmutter. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Simulations were also performed on the national HPE Apollo Hawk at the High Performance Computing Center Stuttgart (HLRS). The authors acknowledge HLRS for funding this project by providing access to the supercomputer HPE Apollo Hawk under the grant numbers INTRHYGUE/44215 and MAGNETIST/44288. Simulations were also performed on SuperMUC_NG at the Leibniz-Rechenzentrum (LRZ) Munich. The authors acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at LRZ (allocations pn67xo, pn76li, pn68wi and pn36jo). Postprocessing and development run were performed on the ARA cluster at Friedrich Schiller University Jena. The ARA cluster is funded in part by DFG grants INST 275/334-1 FUGG and INST 275/363-1 FUGG, and ERC Starting Grant, grant agreement no. BinGraSp-714626.